Grade 12Magical Mathematics[Interesting Approach]

WHAT SHOULD BE THE MISSING DIGIT ? EXPLAIN YOUR ANSWER BY GIVING REASONS.45,41,50, ? ,25,29

Subham Dasgupta
10 Years agoGrade 12

2 Answers

Kanika
10 Years ago
The series will be 45, 41, ?, 50, 25, 29
42 + 52= 41
42 + 12= 17
12 + 72= 50
52 + 02= 25
22 + 52= 29

So, in place of ?, the number will be 17
manmath
10 Years ago
 
= 292 + 52= 2522 + 02= 5052 + 72= 1712 + 12= 4142 + 52 The series will be 45, 41, ?, 50, 25, 294
so number is 17(seventeen)